Operations Report – April 4th

SOL: 7

Name of person filing report: Amanda Manget

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on the generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): 9

Solar – SOC 80% @ 1900 (before David turned on generator for the night)

Diesel Reading – 45%

Propane Reading – 51%

Ethanol Free Gasoline – 2 gallons

Water (auxiliary tank) – 0 gallons (Note: Not in Use)

Water (static tank) – 152.13

Auxiliary to Static tank transfer – No

Gallons transferred: 0

Water in GreenHab – 198 gallons

Static to Loft Pump used – Yes

Water Meter: 0142453.7 gallons @ 1730

Toilet tank emptied: No

Deimos rover used: No, not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: No

Hours: 86

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Opportunity rover used: No

Hours: 56.2

Beginning charge: 100%

Ending charge: 72%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 87.6

Beginning charge: 100%

Ending charge: 55%

Currently charging: Yes

Notes on Rovers: Our EVA#7 Team confirmed the battery charge on the Curiosity and Opportunity Rovers was 100%. GHO drove the Opportunity rover with COM as the lead vehicle for the EVA. At 82% charge, the Opportunity engine stopped at 0517574 4256201 and could not be restarted. After waiting approximately 40 minutes, the Opportunity rover was restarted and continued until 0518227 4250731 before stopping again at 72% charge. Our EVA Team was able to reach David who came out with an ATV to tow the Opportunity rover back to the Hab.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) Honda

Reason for use: Towing Opportunity

Oil Added? No

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 1

Notes on ATVs: None

HabCar used and why, where? Not Used

CrewCar used and why, where? Not Used

CrewCar odometer reading: 175109 miles

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: EV suits normal, all batteries read > 12.5 volts

Summary of Hab operations: Checked SOC of electrical system as follows:

1000 SOC 100%, generator turned off by David

1600 SOC 91%, generator off

1718 SOC 87%, generator off

1900 SOC 80%, generator turned on by David

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: None

EVA Report – April 4th

EVA #7

Author: Julielynn Wong

Purpose of EVA: Geologic exploration and drone flying

Start time: 1430

End time: 1630

Narrative: Our EVA Team confirmed the battery charge on the Curiosity and Opportunity Rovers was 100%. GHO drove the Opportunity rover with COM as the lead vehicle for the EVA. At 82% charge, the Opportunity engine stopped at 0517574 4256201 and could not be restarted. After waiting approximately 40 minutes, the Opportunity rover was restarted and continued until 0518227 4250731 before stopping again at 72% charge. Our EVA Team was able to reach David who came out with an ATV to tow the Opportunity rover back to the Hab.

Destination: Gray Moon (not reached due to rover issues)

Coordinates (use UTM NAD27 CONUS): 0517574 4256201; 0518227 4250731

Participants: Julielynn Wong (COM), Dean Jin (XO), Erika Rydberg (GHO)

Road(s) and routes per MDRS Map: Cow Dung Road 0110, to Rd #1572, to Rd #1575 (Avoid Copernicus Hwy 1574)

Mode of travel: Driving Curiosity + Opportunity rovers

EVA Report – April 4th

EVA # 6

Author: David Kim

Purpose of EVA: Geologic exploration

Start time: 1030h

End time: 1200h

Narrative: Walked up the hill behind the Hab to Hab View North. Further walked to Hab ridge areas and explored local geology and paleontology. No issues encountered.

Destination: Hab View North

Coordinates: x518000, y4251000

Participants: Amanda Manget (ENG), David Kim (HSO)

Road(s) and routes per MDRS Map: n/a

Mode of travel: Walked

Astronomy Report – April 4th

Name: Amanda Manget

Crew: 208
Date: 04-APR-2019

MDRS ROBOTIC OBSERVATORY

Robotic Telescope Requested: MDRS-WF

Objects to be Imaged this Evening (if weather permits):

Submitted March 27th: Pinwheel (active 7/10 images taken)

Submitted April 1st: Crab Nebula, Dumbbell Nebula

Objects Already Imaged: Pinwheel with MDRS-14, Whirlpool Galaxy with MDRS-WF

Problems Encountered: Only a portion of images were taken of the Crab Nebula. If weather permits, more images will be taken tonight.

MUSK OBSERVATORY

Solar Features Observed: N/A

Problems Encountered: N/A

Journalist Report – April 4th

Date: April 4, 2019

Author: Erika Rydberg

Getting into our Groove on Mars

Food is a central part of our day and that is the case, no matter which planet you find yourself on. Our routine helps guide journey here on Mars, it’s part of the reason why I like to include it in each day’s journalist report. We started the day with cereal and fresh oatmeal. This morning David and Amanda left that for a walking EVA to explore the local fossil beds and discovered many, many seashell fossils close by. We had a large print of a map of MDRS finish on one of our 3D printers, and also printed several useful items today on our other printer. We are also using glow-in-the-dark biodegradable filament to 3D print miniature scale models of MDRS structures designed during our mission to place on our 3D printed landscapes. We performed our daily duties through much of the afternoon and continued to perfect our research. We had a lunch of herbed couscous and smoothies and then, Dean, Julielynn and I left for a second rover EVA mission. Amanda and David stayed behind and continued their journeys on our bike-powered 3D printing project, both putting in an hour riding to charge our battery.

Now that we’re nearly done with our mission, we’re starting to wrap up some of our projects and research. With 3D printing, we sometimes think if only there was more time here on the Red Planet, we would be able to accomplish so much more. For those 3D printer enthusiasts out there, you may know a lot of working with these machines is also about patience and waiting. In the limited time we have we’ve been able to gather quite a bit of knowledge in our processes, projects, and research. We hope you’ve enjoyed taking the journey with us, we have only a few days left, until we blast off and return to our home planet.

GreenHab Report – April 4th

GreenHab Officer: Erika Rydberg

Environmental control:

Ambient, door open from 08:30 to 1930.

Shade cloth (80%) on

Average temperatures:

Average: 24.47

Low: 19

High: 32.8 C

Hours of supplemental light: 0

Daily water usage for crops: 5 gallons

Water in Blue Tank— 198 gallons

Time(s) of watering for crops: 8:30 and 13:30

Changes to crops: None

Narrative: Turned on fan to setting 2 to try and keep heat down, made call to David to inform of high temperature, watered plants as instructed at the hottest part of the day.

Harvest: None

Support/supplies needed: None

Sol Summary – April 4th

Sol: 7

Summary Title: Godspeed, little taters! (in sim)

Author’s name: Julielynn Wong

Mission Status: Nominal

Sol Activity Summary:

Last PM, Crew 208 used our bike-charged battery to power our 3D printer to use biodegradable, food-safe plastic to make a customized mallet finger splint design that was originally printed on the ISS.

Today, Crew 208 continued 3D printing in the Hab, including making a wonderful 3D printed drone map of MDRS prepared by GHO!

ENG + HSO observed many fossil specimens during their walking EVA this AM.

Crew ENG + HSO each exercised for 1 hour to generate power for our bike-powered 3D printing study.

Crew 208 broke sim this PM to charge our electric vehicle for the return journey home.

Look Ahead Plan: Continue 3D printing in Hab and data collection for the bike-powered and thoracentesis studies this PM.

Anomalies in work: None

Weather: Sunny with winds

Crew Physical Status: Nominal

EVA: #6 – walking; #7 – Opportunity + Curiousity rovers

Reports to be filed: Operations, Greenhab, Journalist, EVA Report

Support Requested: None